OpenStudy (anonymous):

plz help me Solve for x: 3 over 4 x + 5 over 8 = 4x x = __________________ Write your answer as a fraction in simplest form. Use the "/" symbol for the fraction bar.

OpenStudy (xapproachesinfinity):

\[\frac{ 3 }{\frac{ 4x+5 }{ 8 }}=4x\]

OpenStudy (xapproachesinfinity):

is that what you wrote?